Mary leaves San Francisco and travels toward Los Angeles at 50 mi/h. An hour later, Phil leaves Los Angeles and travels toward San Francisco at 60 mi/h. If the two cities are 380 miles apart, how many hours will it take for Mary to meet Phil?
Answer by ankor@dixie-net.com(22740)

Mary leaves San Francisco and travels toward Los Angeles at 50 mi/h. An hour later, Phil leaves Los Angeles and travels toward San Francisco at 60 mi/h. If the two cities are 380 miles apart, how many hours will it take for Mary to meet Phil?
:
Let t = Mary's time on the road
:
(t-1) = Phil's time on the road; (He left 1 hr later)
:
When they meet, they will have traveled a total 380 miles
:
Write a distance equation; distance = speed * time
:
Mary's distance + Phil's distance = 380
50t + 60(t-1) = 380
:
50t + 60t - 60 = 380
:
110t = 360 + 60
:
t = 440/110
:
t = 4 hrs is Mary's time on the road
:
:
Check solution: (Phils time on the road = 3 hrs
(50*4) + (60*3) =
200 + 180 = 380 mi